OpenStudy (anonymous):

(3x+4)(3x-4) write the steps out for me

OpenStudy (anonymous):

Use this formula: \[\huge \color{green}{(a+b)(a-b) = a^2 - b^2}\] Here a = 3x and b = 4 Can you do it or not??

OpenStudy (saifoo.khan):

HINT: Differences of squares.